Directrix (directrix):

Mary is younger than her husband, but their ages contain the same two digits. The sum of their ages is 11 times the difference of their ages. How old is Mary?

OpenStudy (hoblos):

45

OpenStudy (hoblos):

mary is 45 her husband is 54 45+54 =99 54-45 = 9 45+54 = 11(54-45)

Directrix (directrix):

Correct.

OpenStudy (anonymous):

how did you solve that?

OpenStudy (hoblos):

mary's age =x her husband's age =y x+y = 11(y-x) x+y = 11y-11x 12x=10y 6x=5y since x&y are integers y is divisible by 6 and x is divisible by 5 so the one's digit in x is 5 this means that the ten's digit in y is 5 and y must be divisible by 6 so y=54 thus x=45